A ball is thrown upward with an initial velocity of 15 m/s at an angle of 60.0° above the horizontal. Use energy conservation to find the ball’s greatest height above the ground.


1
Expert's answer
2021-10-28T08:54:31-0400

"mv^2_y/2=mgh\to h=v^2\sin^2\alpha/(2g)" . Answer